The word problem is:

Boats A and B leave the same point at the same time at right angles. B travels 7 km/h slower than A. After 4 hours they are 68 km apart. Find the speed of each boat.
That is exactly how it is written in the book and I can't figure it out. Can you please help me? If so thank you.

Answer by ankor@dixie-net.com(22740)

Boats A and B leave the same point at the same time at right angles.
B travels 7 km/h slower than A. After 4 hours they are 68 km apart. Find the speed of each boat.
:
Let x = boat A's speed
then
(x-7) = boat B's speed
:
Dist = speed times time
A's dist: 4x
B's dist: 4(x-7) = (4x-28)
:
(4x)^2 + (4x-28)^2 = 68^2
16x^2 + 16x^2 - 224x + 784 = 4624
32x^2 - 224x - 3840 = 0
Simplify, divide by 32
x^2 - 7x - 120 = 0
Factors to
(x-15(x+9)
positive solution
x = 15 is boat A's speed
you can find boat b's speed